OpenStudy (anonymous):

Find the exact value of the expression. cos(arcsin1/2+arccotroot 3/3)

hartnn (hartnn):

arcsin 1/2= ? can u evaluate that?

OpenStudy (anonymous):

means \[\sin^{-1} (1/2) \]

OpenStudy (anonymous):

im lost

hartnn (hartnn):

for what angle is the value of sin 1/2 = ?

OpenStudy (anonymous):

60 degrees

hartnn (hartnn):

nopes, sin 60 = root 3/2 sin 30 = 1/2

hartnn (hartnn):

so arcsin 1/2 = 30 degrees, ok ?

OpenStudy (anonymous):

ok

hartnn (hartnn):

and for what angle is cot x = root 3/3 or 1/ root 3 ?

OpenStudy (anonymous):

30?

hartnn (hartnn):

cot 30 - root 3, cot 60 = 1/root 3

hartnn (hartnn):

so u have arccotroot 3/3 = 60 so what is the value of arcsin1/2+arccotroot 3/3 = ?

hartnn (hartnn):

@Zekarias arcsin 1/2 = 60 ?

hartnn (hartnn):

arcsin1/2+arccotroot 3/3 = 30+60= ?

OpenStudy (anonymous):

u lost me

hartnn (hartnn):

u clear with arcsin 1/2 = 30 ?

hartnn (hartnn):

because sin 30 = 1/2

OpenStudy (anonymous):

okay

hartnn (hartnn):

arccotroot 3/3 = 60 because cot 60 =1/ (root 3) = root 3/3

hartnn (hartnn):

so arcsin1/2+arccotroot 3/3 = 30+60= ?

OpenStudy (anonymous):

90 ?

hartnn (hartnn):

yeah, clear till here ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

cos(arcsin1/2+arccotroot 3/3) =cos 90 = ?

OpenStudy (anonymous):

0

hartnn (hartnn):

that is correct and is the final answer :)

OpenStudy (anonymous):

ohh ok thank you

hartnn (hartnn):

welcome ^_^